OpenStudy (anonymous):

sin(pi/4)sin(pi/6) = 1/2 (___(pi/12)- cos (5PI/12) could be most likely sin,cos, tan, etc

OpenStudy (anonymous):

your quetion is not clear enough

OpenStudy (raden):

use the identity : sinA sinB = -1/2 [cos(A+B) - cos(A-B)] = 1/2 [cos(A-B) - cos(A+B)]

OpenStudy (anonymous):

not sure how that applies

OpenStudy (raden):

what's pi/4 - pi/6

OpenStudy (anonymous):

- something?

OpenStudy (raden):

could u calculate difference between 2 fractions, like 1/2 - 1/3 :)

OpenStudy (anonymous):

1/2 - 1/3 = 1/6 ?

OpenStudy (raden):

sure... now calculate pi/4 - pi/6 = ...

OpenStudy (raden):

if u can determine it, the numerator is as the answer for ur question above :p

OpenStudy (anonymous):

pi/12?

OpenStudy (raden):

yes... for pi/4 + pi/6 = 5pi/12 so, the complete answer be cos(pi/12) - cos(5pi/12)

OpenStudy (anonymous):

ty

OpenStudy (raden):

welcome